# Matchings in hypercubes extend to long cycles

### v1

@@ 4.2. Reverse implication (induction step $d-1\rightarrow d$ for $d\geq 6$)
4 We now argue that if $u^{i}\in V(M)$, then $N^{i}_{1}$ and $u^{i}$ satisfy property (H) in Theorem 8 for $Q^{i}_{1}$. If $M^{i}_{1}$ contains a half-layer $L$, then by (ii) and Lemma 15 (i), $L$ is the only half-layer in $M^{i}_{1}$, and by rule (1’), the edge $u^{i}u^{iM}$ is one of its edges, so $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ contains no half-layers of $Q^{i}_{1}$ and property (H) is trivially satisfied. On the other hand, if $M^{i}_{1}$ contains no half-layers, then $N^{i}_{1}$ may only contain a half-layer $L$ if $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}\cup\{u^{M}u^{iM}\}$, which occurs if and only if $u^{M},u^{iM}\in V(Q^{i}_{1})$, and then $L^{\prime}:=L\setminus\{u^{M}u^{iM}\}$ is a near half-layer in $M^{i}_{1}$; see Figure 7 (b). Let $j\in[d]\setminus\{i\}$ be the direction of $L^{\prime}$, which is also the direction of the edge $u^{M}u^{iM}$. By Lemma 15 (i), $M^{i}_{1}$ contains no near half-layers of $Q^{i}_{1}$ in a direction different from $j$, so only $L^{\prime}$ and possibly a second near half-layer in direction $j$ qualify for application of rule (1’). However, the edge $u^{i}u^{iM}$ removed by rule (1’) also has the direction $j$. This is a contradiction, however, as the edges $u^{i}u^{iM}$ and $u^{M}u^{iM}$ have different directions (as $u^{i}\neq u^{M}$).
3 Note that the cycle $C_{0}$ obtained from applying Theorem 8 to extend $N^{i}_{0}\cup P_{0}$ avoids the vertex $z$ by construction. As $N^{i}_{0}\cup P_{0}$ avoids only one other vertex in $Q^{i}_{0}$ apart from $z$, namely the vertex $u$, whose parity is opposite to that of $z$, Lemma 13 implies that $C_{0}$ must also avoid the vertex $u$.
@@
4 We argued before that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We apply Lemma 17 to obtain a perfect matching $P_{0}$ on $K(B_{0})$ so that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. Furthermore, Theorem 5 applies inductively to $K(Q^{i}_{1})$, which does not require any assumptions on $N^{i}_{1}\cup P_{1}$, to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$.
@@
We argued before that $N^{i}_{1}$ and $u^{i}$ satisfy property (H) in Theorem 8 for $Q^{i}_{1}$. Let $I$ be the set of all directions $j\subseteq[d]\setminus\{i\}$ for which there exists a perfect matching $P^{\prime}$ of $K(B_{1})$, where $B_{1}:=V(Q^{i}_{1})\cap V(N^{i}_{-})$ such that $N^{i}_{1}\cup P^{\prime}$ and $u^{i}$ violate property (H), i.e., $N^{i}_{1}\cup P^{\prime}$ contains a half-layer in direction $j$ and all vertices on $u^{i}$’s side of the half-layer are covered. If $j\in I$, then as $u^{i}$ is not covered by $N$, there is exactly one half-layer of $Q^{i}_{1}$ in direction $j$ for which this condition holds. We distinguish three cases depending on the size of $I$.
@@
4 We argued before that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We apply Lemma 17 to obtain a perfect matching $P_{0}$ on $K(B_{0})$ so that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. The definition of $I$ ensures that Theorem 8 applies inductively to $K(Q^{i}_{1})$ (regardless of $C_{0}$) to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$ and avoids $u^{i}$.
@@
4 We argued before that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. Consequently, $N^{i}_{0}\cup\{x^{N}x^{jN}\}$ contains no half-layers of $Q^{i}_{0}$. Thus, we apply Lemma 17 to $N^{i}_{0}\cup\{x^{N}x^{jN}\}$ to obtain a perfect matching $P_{0}$ of $K(Q^{i}_{0})$ that includes the edge $x^{N}x^{jN}$ and contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. As $x^{N}x^{jN}\in P_{0}$, we have $xx^{j}\notin P_{1}$ (regardless of $C_{0}$), and so Theorem 8 applies inductively to $K(Q^{i}_{1})$ to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$ and avoids $u^{i}$.
@@
4 As $N^{i}_{0}$ contains no 2-near half-layers of $Q^{i}_{0}$, we have that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$, so Theorem 8 applies inductively to $K(Q^{i}_{0})$. As $x^{N}x^{jN}\in P_{0}$, we have $xx^{j}\notin P_{1}$ (regardless of $C_{0}$), so Theorem 8 applies inductively to $K(Q^{i}_{1})$.
@@
For any $j\in I$, let $L_{j}$ be the half-layer of $Q^{i}_{1}$ in direction $j$ that is contained in $N^{i}_{1}\cup P^{\prime}$ for some perfect matching $P^{\prime}$ on $K(B_{1})$ (recall that $L_{j}$ is unique as $u^{i}\notin L_{j}$). For any two half-layers $L_{j},L_{k}$, $j,k\in I$, $j\neq k$, Lemma 14 (i) yields that the union $L_{j}\cup L_{k}$ is a collection of paths of length 2. Consequently, if an edge from $L_{j}$ did belong to $N^{i}_{1}$, then its neighboring edge from $L_{k}$ would not belong to $N^{i}_{1}$, but at the same time their common end vertex would not be in $B_{1}$, contradicting the fact that $k\in I$. It follows that $V(L_{j})\subseteq B_{1}$ for all $j\in I$, in other words, none of the edges in any of the half-layers $L_{j}$ belongs to $N^{i}_{1}$. By Lemma 14 (ii), the union $\bigcup_{j\in I}L_{j}$ avoids exactly $2^{d-1-|I|}$ vertices of $Q^{i}_{1}$ and therefore $|N^{i}_{0}|\leq 2^{d-2-|I|}$ and $|N^{i}_{-}|\geq 2^{d-1}-2^{d-1-|I|}$. For every ‘dangerous’ direction $j\in I$ we will pick an edge $xx^{j}\in L_{j}$, and we add the edge $x^{N}x^{jN}$ (which lies in $Q^{i}_{0}$ by the arguments from before) to the matching $P_{0}$, which ensures that $xx^{j}\notin P_{1}$. Specifically, we define $v:=u^{i}$ and we consider the edge $e(j):=v^{p(j)}(v^{p(j)})^{j}\in L_{j}$ for every $j\in[d]$, where $p(j):=j-1\pmod{d}$. Note that the edges $e(j)$, $j\in[d]$, are independent, i.e., no two of them share an end vertex, so for every $j\in I$ we let $xx^{j}:=e(j)$, and we add the edge $x^{N}x^{jN}$ to $P_{0}$. As $|N^{i}_{0}|+|I|\leq 2^{d-2-|I|}+|I|<2^{d-3}$, this does not create half-layers in $N^{i}_{0}\cup P_{0}$. The matching $P_{0}$ on $K(B_{0})$ is completed by applying Lemma 17, using that $|N^{i}_{-}|-2|I|\geq 2^{d-1}-2^{d-1-|I|}-2|I|\geq 4$.
@@ 4.3. Reverse implication (induction basis $d=5$)
It remains to settle the base case $d=5$ in Theorem 8, which we do with computer assistance. In this section we describe the underlying theoretical considerations. Our verification program in C++ is available for download from the authors’ websites [www]. The program is compiled using GCC 12.2.0 on Debian 12, and it runs as a single thread. The reported running times are obtained on an AMD Ryzen 9 7900X3D, 4.4 GHz, 64 GB RAM desktop computer.
@@ 4.3.1. $M$ contains a $z$-dangerous (near) half-layer
If $M$ contains a $z$-dangerous (near) half-layer, then the computer verification proceeds as follows: We generate all matchings of $K(Q_{5})$ that contain a $z$-dangerous near half-layer in some direction $i$ (w.l.o.g. we can assume $i=1$), and for each of them that satisfies property (H), we verify that it can be extended to a Hamilton cycle. We reduce the number of test cases by considering the group of automorphisms generated by permutations of all directions except $i$, and we only consider matchings that are non-isomorphic under this group action. The running time for completing the verifications in this case is about 3 minutes.
@@ 4.3.2. $M$ does not contain $z$-dangerous (near) half-layers
The following arguments are illustrated in Figure 9. We define $A_{1}:=V(Q^{i}_{1})\cap V(M^{i}_{-})$. For some direction $j\in[5]$, $j\neq i$, we define $Q_{1,0}=(Q^{i}_{1})^{j}_{0}$ and $Q_{1,1}=(Q^{i}_{1})^{j}_{1}$. We also define $Q_{1,0}^{\prime}:=Q_{5}\setminus Q_{1,0}$ and $Q_{1,1}^{\prime}:=Q_{5}\setminus Q_{1,1}$. Let $M_{1,0}^{\prime}$ and $M_{1,1}^{\prime}$ be the set of edges of $M$ having both end-vertices in $Q_{1,0}^{\prime}$ and $Q_{1,1}^{\prime}$, respectively. Furthermore, let $A_{1,0}^{\prime}$ and $A_{1,1}^{\prime}$ be the vertices of $Q_{1,0}^{\prime}$ and $Q_{1,1}^{\prime}$ covered by edges of $M$ having the other end-vertex in $Q_{1,0}$ and $Q_{1,1}$, respectively. We write $L^{j}_{0i}$ for the half-layer in $Q^{j}_{0}$ in direction $i$ for which the end vertices whose $i$th bit equals 0 are odd. This is precisely the $z$-dangerous half-layer of $Q^{j}_{0}$ in direction $i$. Similarly, we write $L^{j}_{1i}$ for the half-layer in $Q^{j}_{1}$ in direction $i$ for which the end vertices whose $i$th bit equals 0 are odd. We define :=$\mathcal{L}:=\{L^{j}_{0i},L^{j}_{1i}\mid j\in[d]\setminus\{i\}\}$.
We aim to choose the direction $j\in[5]$, $j\neq i$, so that one of the following two statements is satisfied:
@@
We first consider the case that $M^{i}_{-}$ contains no quad-layers of $Q_{5}$ from the set $\mathcal{L}$. We choose two vertices $x,y\in A_{1}$ and if possible choose them so that $xx^{i},yy^{i}\in M$. We then choose the direction $j$ so that it separates $x$ and $y$, i.e., one is contained in $Q^{j}_{0}$ and the other one in $Q^{j}_{1}$, implying that $|A_{1,0}^{\prime}|>0$ and $|A_{1,1}^{\prime}|>0$. Since $|A_{1}|=|M^{i}_{-}|$ is odd, $|A_{1,0}^{\prime}|$ and $|A_{1,1}^{\prime}|$ have different parity, and so one of them is even and $\geq 2$. By symmetry, we assume w.l.o.g. that $|A_{1,0}^{\prime}|\geq 2$ is even and that $x\in Q_{1,0}$; see Figure 9 (a). We define $B:=A_{1,0}^{\prime}\cap V(Q^{i}_{0})$ and $C:=A_{1,0}^{\prime}\setminus B$. Note that we have $x^{M}\in B$. The matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ is obtained as follows. If $|B|\geq 4$, then we apply Lemma 17 to match all vertices in $B$, except the vertex $x^{M}$ if $B$ is odd, such that no near half-layers of $Q^{i}_{0}$ are created. If $|B|\in\{2,3\}$, then we match one pair of vertices from $B$, except the vertex $x^{M}$ if $|B|=3$, which does not create half-layers of $Q^{i}_{0}$. If $|B|=|\{x^{M}\}|=1$ or the vertex $x^{M}$ from $B$ is still unmatched from the previous cases, we match it arbitrarily to one of the vertices in $C$, and the remaining vertices in $C$ we match arbitrarily among each other. Note that adding an edge from $x^{M}\in B$ to some vertex $c\in C$ cannot create the half-layer $L^{j}_{1i}$, as this would imply $(x^{M})^{i}\in C$ and therefore $(x^{M})^{i}\neq x$ and $xx^{i}\notin M$, but since 3 edges of $L^{j}_{1i}$ are in $M^{i}_{-}$, this contradicts the choice of $x$.
@@
It remains to consider the case that $M^{i}_{-}$ contains a quad-layer $L\in\mathcal{L}$ of $Q_{5}$, which has $42^{d-3}=$ 4 edges. As $M$ does not contain $z$-dangerous (near) half-layers of $Q_{5}$, there are at least two edges $xx^{i}$ and $yy^{i}$ with $x,y\in V(Q^{i}_{1})$ of the $z$-dangerous half-layer in direction $i$ that are not contained in $M^{i}_{-}$; see Figure 9 (c). We choose the direction $j$ so that it separates $x$ and $y$. As before, we obtain that one of $|A_{1,0}^{\prime}|$ and $|A_{1,1}^{\prime}|$ is even and $\geq 2$, and by symmetry we may assume w.l.o.g. that $|A_{1,0}^{\prime}|\geq 2$ is even and that $x\in Q_{1,0}$. As $yy^{i}\in L^{j}_{i1}$ but $yy^{i}\notin M^{i}_{-}$, the matching $M_{1,0}^{\prime}$ does not contain the half-layer $L^{j}_{i1}$ of $Q^{j}_{1}$. Furthermore, as $|L|=4$, at least one of the edges $zz^{i}\in L$ satisfies $z\in Q_{1,0}$. We define $B:=A_{1,0}^{\prime}\cap V(Q^{i}_{0})$ and $C:=A_{1,0}^{\prime}\setminus B$. Note that we have $z^{i}\in B$. The matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ is obtained is follows. If $y^{i}\in B$ and $y\in C$, which happens if $y^{iM},y^{M}\in V(Q_{1,0})$, then we first match $z^{i}$ to $y$ and define $B:=B\setminus\{y^{i}\}$ and $C:=C\setminus\{y\}$. If $|B|\geq 4$, then we apply Lemma 17 to match all vertices in $B$, except one if $B$ is odd, such that no near half-layers of $Q^{i}_{0}$ are created. If $|B|\in\{2,3\}$, then we match one pair of vertices from $B$, except one vertex if $|B|=3$, which does not create half-layers of $Q^{i}_{0}$. If $|B|=1$ or one vertex from $B$ is still unmatched from the previous cases, we match it arbitrarily to one of the vertices in $C$, and the remaining vertices in $C$ we match arbitrarily among each other. Note that $yy^{i}\notin P_{1,0}^{\prime}$, and therefore $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ does not contain the half-layer $L^{j}_{i1}$.

### v2

@@ 4.2. Reverse implication (induction step $d-1\rightarrow d$ for $d\geq 6$)
5 Claim 5b: If $u^{i}\in V(M)$, then $N^{i}_{1}$ contains no half-layers of $Q^{i}_{1}$. It follows in particular that $N^{i}_{1}$ and $u^{i}$ satisfy property (H) in Theorem 8 for $Q^{i}_{1}$ (recall from the proof outline that this property need not be checked if $u^{i}\notin V(M)$). To prove the claim, note that if $M^{i}_{1}$ contains a half-layer $L$ of $Q^{i}_{1}$, then by (ii) and Lemma 15 (i), $L$ is the only half-layer in $M^{i}_{1}$ and all near half-layers must be contained in $L$. By rule (1’), the edge $u^{i}u^{iM}$ is an edge from $L$, so in both of the possible cases $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ or $N^{i}_{1}=(M^{i}_{1}\setminus\{u^{i}u^{iM}\})\cup\{u^{M}u^{iM}\}$ the resulting set $N^{i}_{1}$ contains no half-layers of $Q^{i}_{1}$, and we are done. On the other hand, if $M^{i}_{1}$ contains no half-layers, then $N^{i}_{1}$ may only contain a half-layer $L$ if $N^{i}_{1}=(M^{i}_{1}\setminus\{u^{i}u^{iM}\})\cup\{u^{M}u^{iM}\}$, which occurs if and only if $u^{M},u^{iM}\in V(Q^{i}_{1})$, and then $L^{\prime}:=L\setminus\{u^{M}u^{iM}\}$ is a near half-layer in $M^{i}_{1}$; see Figure 7 (b). Let $j\in[d]\setminus\{i\}$ be the direction of $L^{\prime}$, which is also the direction of the edge $u^{M}u^{iM}$. By Lemma 15 (i), $M^{i}_{1}$ contains no near half-layers of $Q^{i}_{1}$ in a direction different from $j$, so only $L^{\prime}$ and possibly a second near half-layer in direction $j$ qualify for application of rule (1’). However, the edge $u^{i}u^{iM}$ removed by rule (1’) also has the direction $j$. This is a contradiction, however, as the edges $u^{i}u^{iM}$ and $u^{M}u^{iM}$ have different directions (as $u^{i}\neq u^{M}$), proving the claim.
3 $C_{0}$ avoids $u$. Note that the cycle $C_{0}$ obtained from applying Theorem 8 to extend $N^{i}_{0}\cup P_{0}$ avoids the vertex $z$ by construction. As $N^{i}_{0}\cup P_{0}$ avoids only one other vertex in $Q^{i}_{0}$ apart from $z$, namely the vertex $u$ (recall that $X_{0}=\emptyset$), whose parity is opposite to that of $z$, Lemma 13 implies that $C_{0}$ must also avoid the vertex $u$.
@@
4+5 We argued before (Claims 4a and 4b) that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We apply Lemma 17 to obtain a perfect matching $P_{0}$ on $K(B_{0})$ so that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. Furthermore, Theorem 5 applies inductively to $K(Q^{i}_{1})$, which does not require any assumptions on $N^{i}_{1}\cup P_{1}$, to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$.
@@
Let $I$ be the set of all directions $j\in[d]\setminus\{i\}$ for which there exists a perfect matching $P^{\prime}$ of $K(B_{1})$, where $B_{1}:=V(Q^{i}_{1})\cap V(N^{i}_{-})$ such that $N^{i}_{1}\cup P^{\prime}$ and $u^{i}$ violate property (H), i.e., $N^{i}_{1}\cup P^{\prime}$ contains a half-layer in direction $j$ and all vertices on $u^{i}$’s side of the half-layer other than $u^{i}$ itself are covered. If $j\in I$, then as $u^{i}$ is not covered by $N$, there is exactly one half-layer $L$ of $Q^{i}_{1}$ in direction $j$ for which this condition holds. Furthermore, at least one edge of $L$ must come from the matching $P^{\prime}$ and not from $N^{i}_{1}$. This follows from our earlier argument (Claims 5a and 5b) that if $N^{i}_{1}$ contains a half-layer of $Q^{i}_{1}$, then there is a vertex on $u^{i}$’s side of the half-layer that is avoided by $M$ (because this forces $I=\emptyset$).
We distinguish three cases depending on the size of $I$.
@@
4+5 We argued before (Claims 4a and 4b) that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We apply Lemma 17 to obtain a perfect matching $P_{0}$ on $K(B_{0})$ so that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. The definition of $I$ ensures that Theorem 8 applies inductively to $K(Q^{i}_{1})$ (regardless of $C_{0}$) to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$ and avoids $u^{i}$.
@@
4+5 We argued before (Claims 4a and 4b) that $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. Consequently, $N^{i}_{0}\cup\{x^{N}x^{jN}\}$ contains no half-layers of $Q^{i}_{0}$. Thus, we apply Lemma 17 to $N^{i}_{0}\cup\{x^{N}x^{jN}\}$ to obtain a perfect matching $P_{0}$ of $K(B_{0})$ that includes the edge $x^{N}x^{jN}$ such that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$. It follows that Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. As $x^{N}x^{jN}\in P_{0}$, we have $xx^{j}\notin P_{1}$ (regardless of $C_{0}$), and so Theorem 8 applies inductively to $K(Q^{i}_{1})$ to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$ and avoids $u^{i}$.
@@
4+5 As $N^{i}_{0}$ contains no 2-near half-layers of $Q^{i}_{0}$, we have that $N^{i}_{0}\cup P_{0}$ contains no half-layers of $Q^{i}_{0}$, so Theorem 8 applies inductively to $K(Q^{i}_{0})$ to obtain a cycle $C_{0}$ that extends $N^{i}_{0}\cup P_{0}$ and avoids $z$. As $x^{N}x^{jN}\in P_{0}$, we have $xx^{j}\notin P_{1}$ (regardless of $C_{0}$), so Theorem 8 applies inductively to $K(Q^{i}_{1})$ to obtain a cycle $C_{1}$ that extends $N^{i}_{1}\cup P_{1}$ and avoids $u^{i}$.
@@
For any $j\in I$, let $L_{j}$ be the half-layer of $Q^{i}_{1}$ in direction $j$ that is contained in $N^{i}_{1}\cup P^{\prime}$ for some perfect matching $P^{\prime}$ on $K(B_{1})$ (recall that $L_{j}$ is unique as $u^{i}\notin L_{j}$). For any two half-layers $L_{j},L_{k}$, $j,k\in I$, $j\neq k$, Lemma 14 (i) yields that the union $L_{j}\cup L_{k}$ is a collection of paths of length 2. Consequently, if an edge from $L_{j}$ did belong to $N^{i}_{1}$, then its neighboring edge from $L_{k}$ would not belong to $N^{i}_{1}$, but at the same time their common end vertex would not be in $B_{1}$, contradicting the fact that $k\in I$. It follows that $V(L_{j})\subseteq B_{1}$ for all $j\in I$, in other words, none of the edges in any of the half-layers $L_{j}$ belongs to $N^{i}_{1}$. By Lemma 14 (ii), the union $\bigcup_{j\in I}L_{j}$ avoids exactly $2^{d-1-|I|}$ vertices of $Q^{i}_{1}$ and therefore $|N^{i}_{-}|\geq 2^{d-1}-2^{d-1-|I|}$, which implies $|N^{i}_{0}|\leq(2^{d-1}-|N^{i}_{-}|)/2\leq 2^{d-2-|I|}$. For every ‘dangerous’ direction $j\in I$ we will pick an edge $xx^{j}\in L_{j}$, and we add the edge $x^{N}x^{jN}$ (which lies in $Q^{i}_{0}$ by the arguments from before) to the matching $P_{0}$, which ensures that $xx^{j}\notin P_{1}$. Specifically, we define $v:=u^{i}$ and we consider the edge $e(j):=v^{p(j)}(v^{p(j)})^{j}\in L_{j}$ for every $j\in I$, where $p(j):=j-1\pmod{d}$. Note that the edges $e(j)$, $j\in I$, are independent, i.e., no two of them share an end vertex, so for every $j\in I$ we let $xx^{j}:=e(j)$, and we add the edge $x^{N}x^{jN}$ to $P_{0}$. As $|N^{i}_{0}|+|I|\leq 2^{d-2-|I|}+|I|<2^{d-3}$, this does not create half-layers in $N^{i}_{0}\cup P_{0}$. The matching $P_{0}$ on $K(B_{0})$ is completed by applying Lemma 17, using that $|N^{i}_{-}|-2|I|\geq 2^{d-1}-2^{d-1-|I|}-2|I|\geq 4$.
4+5 The argument then continues analogously as in case bii’ before.
@@ 4.3. Reverse implication (induction basis $d=5$)
It remains to settle the base case $d=5$ in Theorem 8, which we do with computer assistance. In this section we describe the underlying theoretical considerations. Our verification program in C++ is available for download from Gitlab [git], and it spans approximately 1400 lines of code. The program is compiled using GCC 12.2.0 on Debian 12, and it runs as a single thread. The reported running times are obtained on an AMD Ryzen 9 7900X3D, 4.4 GHz, 64 GB RAM desktop computer.
Note that $K(Q_{5})$ has $31\cdot 29\cdot 27\cdots 3>10^{17}$ many perfect matchings, considerably more than what could be handled by a naive exhaustive enumeration approach. This explains the following theoretical considerations necessary for being able to perform these verifications in reasonable computing time.
@@ 4.3.1. $M$ contains a $z$-dangerous (near) half-layer
If $M$ contains a $z$-dangerous (near) half-layer, then the computer verification proceeds as follows: We generate all matchings of $K(Q_{5})$ that contain a $z$-dangerous near half-layer in some direction $i$ (w.l.o.g. we can assume $i=1$), and for each of them that satisfies property (H), we verify that it can be extended to a Hamilton cycle that avoids $z$. We reduce the number of test cases by considering the group of automorphisms generated by permutations of all directions except $i$, and we only consider matchings that are non-isomorphic under this group action. The running time for completing the verifications in this case is about 3 minutes.
@@ 4.3.2. $M$ does not contain $z$-dangerous (near) half-layers
The following arguments are illustrated in Figure 9. We define $A_{1}:=V(Q^{i}_{1})\cap V(M^{i}_{-})$. For some direction $j\in[5]$, $j\neq i$, we define $Q_{1,0}:=(Q^{i}_{1})^{j}_{0}$ and $Q_{1,1}:=(Q^{i}_{1})^{j}_{1}$. We also define $Q_{1,0}^{\prime}:=Q_{5}\setminus Q_{1,0}$ and $Q_{1,1}^{\prime}:=Q_{5}\setminus Q_{1,1}$. Let $M_{1,0}:=(M^{i}_{1})^{j}_{0}$ and $M_{1,1}:=(M^{i}_{1})^{j}_{1}$ be the set of edges of $M$ having both end vertices in $Q_{1,0}$ and $Q_{1,1}$, respectively. Similarly, let $M_{1,0}^{\prime}$ and $M_{1,1}^{\prime}$ be the set of edges of $M$ having both end vertices in $Q_{1,0}^{\prime}$ and $Q_{1,1}^{\prime}$, respectively. Furthermore, let $M_{1,0,-}$ and $M_{1,1,-}$ be the edges of $M$ having exactly one end vertex in $Q_{1,0}$ and $Q_{1,1}$, respectively, and let $A_{1,0}^{\prime}$ and $A_{1,1}^{\prime}$ be the vertices of $Q_{1,0}^{\prime}$ and $Q_{1,1}^{\prime}$ covered by edges of $M_{1,0,-}$ and $M_{1,1,-}$. We write $L^{j}_{0i}$ for the half-layer in $Q^{j}_{0}$ in direction $i$ for which the end vertices whose $i$th bit equals 0 are odd. This is precisely the $z$-dangerous half-layer of $Q^{j}_{0}$ in direction $i$. Similarly, we write $L^{j}_{1i}$ for the half-layer in $Q^{j}_{1}$ in direction $i$ for which the end vertices whose $i$th bit equals 0 are odd. We define :=$\mathcal{L}:=\{L^{j}_{0i},L^{j}_{1i}\mid j\in[d]\setminus\{i\}\}$.
We aim to choose the direction $j\in[5]$, $j\neq i$, so that one of the following two conditions is satisfied:
  1. (a)
@@
We first consider the case that $M^{i}_{-}$ contains no quad-layers of $Q_{5}$ from the set $\mathcal{L}$. We choose two vertices $x,y\in A_{1}$ and if possible choose them so that $xx^{i},yy^{i}\in M$. We then choose the direction $j$ so that it separates $x$ and $y$, i.e., one is contained in $Q^{j}_{0}$ and the other one in $Q^{j}_{1}$, implying that $|A_{1,0}^{\prime}|>0$ and $|A_{1,1}^{\prime}|>0$. Since $|A_{1}|=|M^{i}_{-}|$ is odd, $|A_{1,0}^{\prime}|$ and $|A_{1,1}^{\prime}|$ have different parity, and so one of them is even and at least 2. By symmetry, we assume w.l.o.g. that $|A_{1,0}^{\prime}|\geq 2$ is even and that $x\in Q_{1,0}$; see Figure 9 (a). We define $B:=A_{1,0}^{\prime}\cap V(Q^{i}_{0})$ and $C:=A_{1,0}^{\prime}\setminus B$. Note that we have $x^{M}\in B$. The matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ is obtained as follows. If $|B|\geq 4$, then we apply Lemma 17 to match all vertices in $B$, except the vertex $x^{M}$ if $B$ is odd, such that no near half-layers of $Q^{i}_{0}$ are created. If $|B|\in\{2,3\}$, then we match one pair of vertices from $B$, except the vertex $x^{M}$ if $|B|=3$, which does not create half-layers of $Q^{i}_{0}$. If $|B|=|\{x^{M}\}|=1$ or the vertex $x^{M}$ from $B$ is still unmatched from the previous cases, we match it arbitrarily to one of the vertices in $C$, and the remaining vertices in $C$ we match arbitrarily among each other. Note that adding an edge from $x^{M}\in B$ to some vertex $c\in C$ cannot create the half-layer $L^{j}_{1i}$, as this would imply $(x^{M})^{i}\in C$ and therefore $(x^{M})^{i}\neq x$ and $xx^{i}\notin M$, but since 3 edges of $L^{j}_{1i}$ are in $M^{i}_{-}$, this contradicts the choice of $x$.
@@
It remains to consider the case that $M^{i}_{-}$ contains a quad-layer $L\in\mathcal{L}$ of $Q_{5}$, which has $42^{d-3}=$ 4 edges. As $M$ does not contain $z$-dangerous (near) half-layers of $Q_{5}$, there are at least two edges $xx^{i}$ and $yy^{i}$ with $x,y\in V(Q^{i}_{1})$ of the $z$-dangerous half-layer in direction $i$ that are not contained in $M^{i}_{-}$; see Figure 9 (c). We choose the direction $j$ so that it separates $x$ and $y$. As before, we obtain that one of $|A_{1,0}^{\prime}|$ and $|A_{1,1}^{\prime}|$ is even and at least 2, and by symmetry we may assume w.l.o.g. that $|A_{1,0}^{\prime}|\geq 2$ is even and that $x\in Q_{1,0}$. As $yy^{i}\in L^{j}_{1i}$ but $yy^{i}\notin M^{i}_{-}$, the matching $M_{1,0}^{\prime}$ does not contain the half-layer $L^{j}_{1i}$ of $Q^{j}_{1}$. Furthermore, as $|L|=4$, at least one of the edges $zz^{i}\in L$ satisfies $z\in Q_{1,0}$. We define $B:=A_{1,0}^{\prime}\cap V(Q^{i}_{0})$ and $C:=A_{1,0}^{\prime}\setminus B$. Note that we have $z^{i}\in B$. The matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ is obtained is follows. If $y^{i}\in B$ and $y\in C$, which happens if $y^{iM},y^{M}\in V(Q_{1,0})$, then we first match $z^{i}$ to $y$ and define $B:=B\setminus\{y^{i}\}$ and $C:=C\setminus\{y\}$. If $|B|\geq 4$, then we apply Lemma 17 to match all vertices in $B$, except one if $B$ is odd, such that no near half-layers of $Q^{i}_{0}$ are created. If $|B|\in\{2,3\}$, then we match one pair of vertices from $B$, except one vertex if $|B|=3$, which does not create half-layers of $Q^{i}_{0}$. If $|B|=1$ or one vertex from $B$ is still unmatched from the previous cases, we match it arbitrarily to one of the vertices in $C$, and the remaining vertices in $C$ we match arbitrarily among each other. Note that $yy^{i}\notin P_{1,0}^{\prime}$, and therefore $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ does not contain the half-layer $L^{j}_{1i}$.